Graph the given​ functions, f and​ g, in the same rectangular coordinate system. Describe how the graph of g is related to the graph of
f(x)= -x^3
​g(x)= -x^3-2
1
Expert's answer
2020-08-24T18:47:47-0400

"f(x)=-x^3\\
g(x)=-x^3-2\\"

Transformation: shift a graph of y=f(x) 2 units down.

Graphs of f(x) and g(x) in the same rectangular coordinate system.